Armond has two solutions available in the lab, one with 5% alcohol and another with 13% alcohol. How much of each solution should he mix together to obtain 6 gallons of a solution that contains 8% alcohol.

I have no idea where to even begin on this problem...

Answer by stanbon(75887)

Armond has two solutions available in the lab, one with 5% alcohol and another with 13% alcohol. How much of each solution should he mix together to obtain 6 gallons of a solution that contains 8% alcohol.
---------------------------------------
alcohol + alcohol = alcohol
0.06x + 0.13(6-x) = 0.08*6
-------
Multiply thru by 100 to get:
6x + 13*6 - 13x = 8*6
-7x = -5*6
x = 30/7
----
x = 4 2/7 gallons (amt. of 6% solution)
----
6-x = 1 5/7 gallons (amt. of 13% solution)
=================================================
Cheers,
Stan H.
================